a merchant has coffee worth $2 a pound that she wishes to mix with 9 pounds of coffee worth $8 a pound to get a mixture that can be sold for $4 a pound. how many pounds of the $2 coffee should be used?
Answer by mananth(16946)

--- price ---------------- Amount
Coffee I 8 ---------------- 9 lbs
Coffee 2 2 ---------------- x lbs
Mixture 4 ---------------- 9 + x lbs

8 * 9 + 2 x = 4 ( 9 + x )
72 + 2 x = 36 + 4 x
2 x -4 x = -72 + 36
-2 x = -36
/ -2
x = 18 lbs Coffee 2

m.ananth@hotmail.ca